OpenStudy (notamathgenius):

The perimeter of a basketball court is 114 meters and the length is 6 meters longer than twice the width. What are the length and width?

OpenStudy (notamathgenius):

@benjii

OpenStudy (notamathgenius):

@Gokuporter

OpenStudy (wolf1728):

You need 2 equations: A) perimeter = 2*length + 2*width = 114 B) Length = width + 6 neatening that up A) 2L + 2W = 114 B) Length - Width = 6 Can you solve those equations?

OpenStudy (wolf1728):

Multiplying equation B by -2 A) 2L + 2W = 114 B) -2L + 2W = -12 4 W = 102 Width = 25.5 Length = Width + 6 Length = 31.5